Please explain step by step and write clearly. Thank

Solve the IVP 3y 2y 0 y(0)=3 y(0)=-5
Answer #1

3Yty-zyzo yo)=3, ylo) =-5 - A .sgnis 3m²fm-2=0 3m? +3m2m-270 3 m m + 1) - 21 + -0 Bm-2) (m+bro 2, you = cent q e if y los =